# Stable-V2A: Synthesis of Synchronized Sound Effects with Temporal and Semantic Controls

### v2

@@ IV-C Stage 1: RMS-Mapper
The first stage of our model, the RMS-Mapper, is a simple neural network that takes some features extracted from the video as input and generates a temporal guide for the sound synthesis process of Stage 2.
@@
The target is an envelope extracted from a corresponding 10 seconds audio chunk sampled at 16 kHz. This relatively low sample rate helps reduce the temporal resolution gap between audio and video while preserving sufficient detail in the audio track. We compute the RMS with the function implemented in the Librosa Feature module [^1] https://librosa.org/doc/main/generated/librosa.feature.rms.html. Following the nomenclature of Eq. 7, we set $W=512$ and $h=128$, thus obtaining an envelope $\mathbf{r}$ with 1250 time samples. Since we normalize the audio $\mathbf{y}$ in the $[-1,1]$ range, the resulting $\mathbf{r}$ is a curve with values in the range $[0,1]$.
@@
In addition, to mitigate the high-frequency variations which are common in real sound waveforms, we apply a smoothing filter with a kernel size of 15. Our experiments show that incorporating such a filter significantly improves the results, simplifying the classification task that the network performs on the video frames, as can be seen in Fig. 3.
Finally, we map the time-continuous RMS-envelope to 64 distinct classes, by applying $\mu$-law encoding, as reported in Eq. 8, using the function provided in the Librosa library [^2] https://librosa.org/doc/main/generated/librosa.mu_compress.html. We then apply Gaussian label smoothing on the target discretized envelope.
@@
RMS-Mapper takes as input features extracted from the videos. We first convert each video to 30 fps, doubling the frame rate compared to our previous work [30]. This means that a 10 seconds long video is represented by a sequence of 300 consecutive frames.
In addition, we compute the optical flow of each video using the RAFT [68] model. Experimentally, we have noticed that the information contained in the optical flow is crucial for inferring a relationship between the motion in the video and the corresponding sound.
@@
We use TC-CLIP [69] to extract relevant features from the RGB frames and the optical flow. Temporally Contextualized CLIP (TC-CLIP) is a model for video understanding that leverages essential temporal information through global interactions in a spatio-temporal domain within a video. TC-CLIP generates 512 dimensional frame-wise features. We concatenate the RGB and optical flow features along the feature dimension, obtaining an input of size $[\mathrm{video\_frames},1024]$, where in our case $\mathrm{video\_frames}=300$ for Greatest Hits experiments.
@@
Then, we use an efficient model based on the RegNet video encoder [32], that is the same baseline of [54], and we improve it to generate the output of our model. The network is composed of three 1D convolutional layers, a two-layer bidirectional LSTM (Bi-LSTM) [70] and a linear layer that projects the resulting features to the output size required for the classification task. We use the convolutional block as an upsample branch needed to scale features from video temporal resolution to the audio temporal resolution. Each convolutional layer is followed by a batch normalization layer, a ReLU activation function and an Upsample layer. The upsample sizes for the three layers are $[600,1200,1250]$, thus increasing the time dimension of the output from $\mathrm{video\_frames}$ to $\mathrm{RMS\_frames}$, where $\mathrm{RMS\_frames}$ is equal to $1250$. Thanks to this upsample, we do not perform a frame-by-frame classification of the video, but we directly map the continuous RMS envelope.
Figure 4: Block diagram for the proposed RMS-Mapper.
@@
where $\mathbf{r}_{\mathbf{d}}$ is the discretized RMS envelope $\mathbf{r}$, i.e. the envelope after $\mu$-law encoding and Gaussian label smoothing, and $\hat{\mathbf{r}}_{\mathbf{d}}$ is the output of our model. Fig. 4 depicts a block diagram of the proposed RMS-Mapper.
IV-D Stable-Foley
Stable-Foley is based on Stable Audio Open [33], which is an Audio Latent Diffusion Model for generating long-form, variable-length stereo music and sounds at 44.1 kHz using text prompts. Therefore, this model is not specifically trained for V2A tasks, which means that the generated audio does not take into account a time alignment with a specific reference.
We want to use the prior knowledge of such a state-of-the-art model for audio synthesis, but at the same time guide audio generation both semantically and temporally. This means that the audio synthesis model should be controlled by the envelope extracted from the RMS-Mapper, representing our temporal control, and conditionings that capture the semantics and duration of the audio.
@@ IV-D Stable-Foley
Stable-Foley is trained on the same loss on which Stable Audio models are trained. As stated in [61], the final loss is the result of the combination of multiple losses: 1) a reconstruction loss based on a perceptually weighted multiresolution STFT [71]. This loss is computed on both the mid-side (M/S) and the left-right (L/R) representations of stereo audio. To address potential ambiguity in L/R channel placement, the L/R component is weighted at 0.5 relative to the M/S component. 2) an adversarial loss with feature matching incorporated, leveraging five convolutional discriminators [72]. 3) a KL divergence loss term with a weighting factor of $10^{-4}$.
@@ IV-D1 Temporal Control
We use a ControlNet to drive generation using a temporal control and at the same time just fine-tune Stable Audio Open on our specific task, without having to train the model from scratch.
To the best of our knowledge, this is the first attempt to use a ControlNet to control the DiT generation process of Stable Audio with additional control. As ControlNet was originally developed for Diffusion U-Net, we followed what is done in [62] for our implementation of such network for Stable Audio.
@@
The input of the ControlNet needs to be of the same dimension as the input of the diffusion model, i.e. the embedding of the audio used in Stable Audio. To extract this latent representation of the envelope, we use the VAE introduced in [60] for encoding the input to the diffusion model, without needing to update its weights. This VAE downsamples the input stereo audio by a factor of 1024. Specifically, it maps an input signal $\mathbf{y}\in\mathbb{R}^{2\times L}$, where $L$ is the number of samples representing the waveform, and $2$ is the number of channels, to an embedding $\mathbf{y_{c}}\in\mathbb{R}^{64\times\frac{L}{1024}}$.
@@
For the design of Stable-Foley we follow the general architecture of ControlNet: the DiT layers are then frozen, and a trainable copy for each of them is created with two zero-initialized convolutional layers placed before and after the copy. For training, the envelope is generated so that it has the same length in samples as the target audio. The control signal $\mathbf{r}$, representing the RMS envelope, and the input waveform $\mathbf{y}$ are both encoded through the same VAE. For each layer, the encoded control signal $\mathbf{r_{c}}$ is first processed through the initial zero-initialized convolutional layer, then added to the input $\mathbf{y_{c}}$, which is the latent version of $\mathbf{y}$ obtained through the VAE; the resulting signal is subsequently passed through the trainable copy and the second zero-initialized convolutional layer and finally added to the output of the frozen layer, derived from input $\mathbf{y_{c}}$.
@@ IV-D2 Semantic Control
Lots of recent deep learning research has focused on developing versatile audio representations that can generalize effectively across various downstream tasks [73]. Contrastive learning, in particular, gained great popularity for the training of multimodal models. A notable example of this approach is CLAP [35], which aligns embeddings for both audio and text in a shared latent space. We then condition our sound synthesis model on CLAP audio embeddings during training, enabling it to incorporate text-based conditioning as an additional modality at inference time only.
To further refine the semantic and temporal alignment of the final audio, we also condition the generation process with a direct frame representation of the reference video. We do this using the latent representation provided by the CAVP video encoder, introduced in [51]. This video encoder aligns the audio and video modalities, providing useful information about the timing and semantics of the related audio. CAVP takes as input video at 4 fps, so we use this frame rate to pass only RGB frames to it.
These semantic controls are provided via cross-attention layers, as originally proposed for Stable Audio global conditionings.
Conditionings on duration, provided by the hyperparameters $\mathrm{seconds\_start}$ and $\mathrm{seconds\_total}$, indicating the total length of the audio, are fed to the model via prepending it to the input of the model. Prepend conditioning includes also timestep conditioning, indicating the current diffusion timestep. Prepend conditionings are concatenated along the channels dimension to the input.
Figure 5: Block diagram for Stable-Foley. Trainable modules are ControlNet blocks while Stable DiT blocks are frozen.
@@ V-A1 Greatest Hits
The high quality of the samples in this dataset is critical for training V2A models, as very often datasets of video in the wild do not guarantee either good video and audio quality and sufficient audiovisual alignment to allow the models to learn the relationships required to generate audio that is semantically and temporally aligned with the reference video. Altogether, the dataset consists of 977 videos captured both outdoor and indoor. Indoor scenes contain a variety of hard and soft materials, such as metal, plastic, cloth, while the outdoor scenes contain materials that scatter and deform, such as grass, leaves and water. On average, each video contains 48 actions, divided between hitting and scratching. This ensures that each extracted chunk of video, lasting either 10s, contains a sufficient number of hits. We divide the dataset into 732 videos for the training set, 49 for the validation set and 196 for the test set.
@@ V-A2 Walking The Maps
We want to test our model on a case study of interest, among the most common in the sound design of audiovisual works: the generation of footsteps sounds. Datasets properly adapted for V2A tasks do not exist or are not currently publicly available. A dataset for V2A models must consist of high-resolution video and sound design-quality audio.
We therefore decided to build a small dataset with these characteristics in order to test our model in real-world scenarios for Foley synthesis. The audio and video quality of the most modern video games is extremely high, making them the perfect source, in our opinion, to create a V2A dataset. Therefore, we collected clips taken from publicly available YouTube videos of walkthroughs of some famous video games.
A lot of gamers upload videos to YouTube in which they wander around with their animated character in the maps of different video games. These videos are the perfect target for our dataset, in fact the sound of footsteps is clearly audible in many cases and each footstep is strongly characterized, e.g. steps on grass, concrete or wood sound very different from each other as well as steps related to a slow walk or a run. High-definition video is associated with such sounds for more modern video games, allowing for an excellent temporal and semantic relationship between audio and video.
@@
In order to create a first small version of our dataset, called Walking The Maps, we chose 4 video games as targets: Hogwarts Legacy, Zelda Breath of the Wild, Assassin’s Creed: Odyssey, Assassin’s Creed IV Black Flag. For each selected video, we extracted only clips in which the sound of the steps is clearly audible and there are no other possible sound sources in the video that can be related to the target sound.
Thus, our dataset is finally composed of 893 video clips of different lengths. The average duration of the videos is 8.82 seconds, where the shortest video has a duration of 2.04 seconds and the longest one is 72.05 seconds long. Each chunk is saved by reporting the unique ID of the video posted on YouTube, the start second of the chunk related to the full video, and the end instant expressed in seconds, so each video in the dataset will have a name of the type $\mathrm{ID\_start\_end}$.mp4. The audio related to each video was subsequently preprocessed with the AudioSep [74] model using the textual query “footstep sounds” in order to separate as much as possible the footstep sound from any other noise, obtaining as clean a target audio as possible.
Since the shortest video in our dataset is 2 seconds long and our intent is mostly demonstrative, we fine-tuned our Stable-V2A model on 2 seconds chunks of each video in the dataset. The dataset is made publicly available so that it can be used for evaluation of V2A models and to be further extended.
@@ V-B Training and Inference Details
We train RMS-Mapper and Stable-Foley separately. The Stage 1 model is trained on a single 48 GB Nvidia RTX A6000. We use a batch size of 64, training the model for 500 epochs with a learning rate maintained constant at $1\times 10^{-3}$ and Adam as an optimizer with a weight decay set to $1\times 10^{-3}$.
To train Stable-Foley, we use the official Stable Audio Open [^3] https://github.com/Stability-AI/stable-audio-tools repository and the related checkpoint [^4] https://huggingface.co/stabilityai/stable-audio-open-1.0 to initialize the weights of our model. We use 44.1 kHz stereo audio as ground truth for experiments with Greatest Hits. Stable-Foley is trained on a single 48 GB Nvidia RTX A6000 with a batch size of 12 for 20k steps. We use a fixed learning rate of $1\times 10^{-4}$ and AdamW as an optimizer, with parameters set as it is done in Stable Audio Open.
In inference, we use the RMS envelopes predicted though the RMS-Mapper, and after $\mu$-law decoding [^5] https://librosa.org/doc/main/generated/librosa.mu_expand.html, we use interpolation to match the sample rate and employ them as input to the ControlNet of Stable-Foley. The audio synthesis model, then, takes 150 sampling steps to generate the final output, using classifier-free guidance [75] with a scale set to 2.
Regarding the experiments on our Walking The Maps dataset, we fine-tuned our model starting from the best checkpoints obtained after training on Greatest Hits for both RMS-Mapper and Stable-Foley. In order to select chunks with as little background noise and camera movements as possible, we extracted 2 seconds non-overlapping chunks from each video. This implies a change in the temporal dimensions of the data in our model. Other than that, the parameters used for these experiments are the same as those used for Greatest Hits.
In this case, RMS-Mapper is fine-tuned with a decreasing learning rate scheduler, with $\gamma=0.5$, for 2500 epochs. While Stable-Foley is fine-tuned without any change with respect to Greatest Hits experiments.

### v3

@@ IV-C Video Model
The first stage of our model, the video model, is a simple neural network that takes some features extracted from the video as input and generates a temporal guide for the sound synthesis process of the second stage.
@@
The target is an envelope extracted from a corresponding 10-second audio chunk sampled at 16 kHz. This relatively low sample rate helps reduce the temporal resolution gap between audio and video while preserving sufficient detail in the audio track. We compute the RMS with the function implemented in the Librosa Feature module [^1] https://librosa.org/doc/main/generated/librosa.feature.rms.html. Following the nomenclature of Eq. 7, we set $W=512$ and $h=128$, thus obtaining an envelope $\mathbf{r}$ with 1250 time samples. Since we normalize the audio $\mathbf{y}$ in the $[-1,1]$ range, the resulting $\mathbf{r}$ is a curve with values in the range $[0,1]$.
@@
In addition, to mitigate the high-frequency variations which are common in real sound waveforms, we apply a smoothing filter with a kernel size of 15. Our experiments show that incorporating such a filter is essential to improve the results, simplifying the classification task that the network performs on the video frames, as can be seen in Fig. 3. Finally, we map the time-continuous RMS-envelope to 64 distinct classes by applying $\mu$-law encoding, using the function provided in the Librosa library [^2] https://librosa.org/doc/main/generated/librosa.mu_compress.html. We then apply Gaussian label smoothing on the target discretized envelope.
@@
the video model takes as input features extracted from the videos. We first convert each video to 30 fps. This means that a 10-second long video is represented by a sequence of 300 consecutive frames.
In addition, we compute the optical flow of each video using the RAFT [58] model. To effectively capture both appearance and motion cues relevant to sound generation, we utilize features extracted by Temporally Contextualized CLIP (TC-CLIP) [59] from both RGB frames and RAFT-computed optical flow. TC-CLIP excels at video understanding by integrating temporal context. Our experiments confirm that combining static frame appearance (RGB) with motion information (optical flow) via TC-CLIP significantly improves the ability of the video model to predict accurate audio envelopes, as motion is often a key indicator of sound events.
@@
TC-CLIP generates 512 dimensional frame-wise features. We concatenate the RGB and optical flow features along the feature dimension, finally obtaining an input of size $[\mathrm{video\_frames},1024]$, where in our case $\mathrm{video\_frames}=300$ for Greatest Hits experiments.
@@
Then, we use an efficient model based on the RegNet video encoder [34], and we improve it to generate the output of our model. The network is composed of three 1D convolutional layers, a two-layer bidirectional LSTM (Bi-LSTM) [60] and a linear layer that projects the resulting features to the output size required for the classification task. We use the convolutional block as an upsample branch needed to scale features from video temporal resolution to the audio temporal resolution. Each convolutional layer is followed by a batch normalization layer, a ReLU activation function and an Upsample layer. The upsample sizes for the three layers are $[600,1200,1250]$, thus increasing the time dimension of the output from $\mathrm{video\_frames}$ to $\mathrm{RMS\_frames}$, where $\mathrm{RMS\_frames}$ is equal to $1250$. Thanks to this upsample, we do not perform a frame-by-frame classification of the video, but we directly map the continuous RMS envelope.
Figure 4: Block diagram for the proposed the video model.
@@
where $\mathbf{r}_{\mathbf{d}}$ is the discretized RMS envelope $\mathbf{r}$, i.e. the envelope after $\mu$-law encoding and Gaussian label smoothing, and $\hat{\mathbf{r}}_{\mathbf{d}}$ is the output of our model. Fig. 4 depicts a block diagram of the proposed the video model.
IV-D Audio Synthesis Model
The audio synthesis model leverages Stable Audio Open [61], which is an Audio Latent Diffusion Model for generating long-form, variable-length stereo music and sounds at 44.1 kHz using text prompts. Therefore, this model is not specifically trained for V2A tasks, which means that the generated audio does not take into account a time alignment with a specific reference.
We want to use the prior knowledge of such a state-of-the-art model for audio synthesis, but at the same time guide audio generation both semantically and temporally. This means that the audio synthesis model should be controlled by the envelope extracted from the the video model, representing our temporal control, and conditionings that capture the semantics and duration of the audio.
@@ IV-D Audio Synthesis Model
The audio synthesis model is trained on the same L2 loss on which Stable Audio models are trained.
@@ IV-D1 Temporal Control
We use a ControlNet to drive generation using a temporal control and at the same time just fine-tune Stable Audio Open on our specific task, without having to train the model from scratch. A key innovation of the audio synthesis model lies in its temporal control mechanism. We adapt the ControlNet framework, originally designed for U-Nets, to operate with Diffusion Transformer (DiT) architecture of Stable Audio, following recent advancements [52]. To our knowledge, this represents the first successful integration of ControlNet with Stable Audio’s DiT to impose explicit, fine-grained temporal control via an external signal like the RMS envelope, allowing us to precisely guide the synthesis process while leveraging the power of the pre-trained Stable Audio model.
@@
The input of the ControlNet needs to be of the same dimension as the input of the diffusion process, i.e. the embedding of the audio used in Stable Audio. To extract this latent representation of the envelope, we use the VAE introduced in [49] for encoding the input to the diffusion model, without needing to update its weights. This VAE downsamples the input stereo audio by a factor of 1024. Specifically, it maps an input signal $\mathbf{y}\in\mathbb{R}^{2\times L}$, where $L$ is the number of samples representing the waveform, and $2$ is the number of channels, to an embedding $\mathbf{y_{c}}\in\mathbb{R}^{64\times\frac{L}{1024}}$.
@@
For the design of the audio synthesis model we follow the general architecture of ControlNet: the DiT layers are then frozen, and a trainable copy for each of them is created with two zero-initialized convolutional layers placed before and after the copy. For training, the envelope is computed so that it has the same length in samples as the target audio. The control signal $\mathbf{r}$, representing the RMS envelope, and the input waveform $\mathbf{y}$ are both encoded through the same VAE. For each layer, the encoded control signal $\mathbf{r_{c}}$ is first processed through the first zero-initialized convolutional layer, then added to the input $\mathbf{y_{c}}$, which is the latent version of $\mathbf{y}$ obtained through the VAE; the resulting signal is subsequently passed through the trainable copy and the second zero-initialized convolutional layer and finally added to the output of the frozen layer, derived from input $\mathbf{y_{c}}$. The visualization of this process is depicted in Fig. 5.
@@ IV-D2 Semantic Control
Several recent deep learning investigations have focused on developing versatile audio representations that can be effectively generalized across various downstream tasks [62]. Contrastive learning, in particular, gained great popularity for the training of multimodal models. A notable example of this approach is CLAP [63], which aligns embeddings for both audio and text in a shared latent space. We then condition our sound synthesis model on CLAP audio embeddings during training, enabling it to incorporate text-based conditioning as an additional modality at inference time only.
To further refine the semantic and temporal alignment of the final audio, we also condition the generation process with a direct frame representation of the reference video. We do so using the embeddings provided by the CAVP video encoder, introduced in [37]. This video encoder aligns the audio and video modalities, providing useful information about the timing and semantics of the related audio. CAVP takes as input video at 4 fps, so we use this frame rate to pass only RGB frames to it. These semantic controls are fed via cross-attention layers, as global conditionings for Stable Audio.
Conditionings on duration, represented by the hyperparameters $\mathrm{seconds\_start}$ and $\mathrm{seconds\_total}$, indicating the total length of the audio, are fed to the model via prepending it to the input of the model. Prepend conditioning also includes timestep conditioning, indicating the current diffusion timestep. These conditionings are concatenated along the channels dimension to the input.
Figure 5: Block diagram for the audio synthesis model. Trainable modules are ControlNet blocks while Stable DiT blocks are frozen.
@@ V-A1 Greatest Hits
The high quality of the samples in this dataset is critical for training V2A models, as very often datasets of video in the wild do not guarantee either good video and audio quality and sufficient audiovisual alignment to allow the models to learn the relationships required to generate audio that is semantically and temporally aligned with the reference video. Altogether, the dataset consists of 977 videos captured both outdoor and indoor. Indoor scenes contain a variety of hard and soft materials, such as metal, plastic, cloth, while the outdoor scenes contain materials that scatter and deform, such as grass, leaves and water. On average, each video contains 48 actions, divided between hitting and scratching. This ensures that each extracted chunk of video, lasting either 10 seconds, contains a sufficient number of hits. We divide the dataset into 732 videos for the training set, 49 for the validation set and 196 for the test set.
@@ V-A2 Walking The Maps
We want to train and test our model on a case study of interest, among the most relevant in the sound design of audiovisual works: the generation of footsteps sounds. Datasets properly adapted for V2A tasks do not exist or are not currently publicly available. A dataset for V2A models must consist of high-resolution video and sound design-quality audio. We therefore decided to build a novel dataset with these characteristics in order to test our model in real-world scenarios for Foley synthesis. The audio and video quality of the most modern video games is extremely high, making them the perfect source, in our opinion, to create a V2A dataset. Then, we collected clips taken from publicly available YouTube videos of walkthroughs of some famous video games.
Many gamers upload videos to YouTube in which they roam with their animated character in the maps of different video games. These videos are the perfect target for our dataset; in fact, the sound of footsteps is clearly audible and each footstep is strongly characterized, e.g. steps on grass, concrete or wood sound completely different from each other as well as steps related to a slow walk or a run. High-definition video is associated with such sounds for more modern video games, allowing for an excellent temporal and semantic relationship between audio and video.
@@
In order to create the current version of our dataset, called Walking The Maps, we chose 4 video games as targets: Hogwarts Legacy, Zelda Breath of the Wild, Assassin’s Creed: Odyssey, Assassin’s Creed IV Black Flag. For each selected video, we extracted only clips in which the sound of the steps is clearly audible and there are no other possible sound sources in the video that can be related to the target sound.
Thus, our dataset is finally composed of 893 video clips of different lengths. The average duration of the videos is 8.82 seconds, where the shortest video has a duration of 2.04 seconds and the longest one is 72.05 seconds long. Each chunk is saved by reporting the unique ID of the video posted on YouTube, the start second of the chunk related to the full video, and the end instant expressed in seconds, so each video in the dataset will have a name of the type $\mathrm{ID\_start\_end}$.mp4. While YouTube game walkthroughs often feature prominent footstep sounds, they can also contain background music, dialogue, or other game sounds. To ensure the dataset provides clean target audio primarily containing footsteps, essential for training robust V2A models, we preprocessed the audio from each clip using the AudioSep [64] source separation model with the query “footstep sounds”.
Since the shortest video in our dataset is 2 seconds long, we fine-tuned our FOL·AI model on 2 seconds chunks of each video in the dataset. The dataset is made publicly available so that it can be used for evaluation of V2A models and to be further extended.
@@ V-B Training and Inference Details
We train the video model and the audio synthesis model separately. The video model is trained on a single 48 GB Nvidia RTX A6000. We use a batch size of 64, training the model for 500 epochs with a learning rate maintained constant at $1\times 10^{-3}$ and Adam as an optimizer with a weight decay set to $1\times 10^{-3}$.
To train the audio synthesis model, we use the official Stable Audio Open [^3] https://github.com/Stability-AI/stable-audio-tools repository and the related checkpoint [^4] https://huggingface.co/stabilityai/stable-audio-open-1.0 to initialize the weights of our model. We use 44.1 kHz stereo audio as ground truth for experiments with Greatest Hits. the audio synthesis model is trained on a single 48 GB Nvidia RTX A6000 with a batch size of 12 for 20k steps. We use a fixed learning rate of $1\times 10^{-4}$ and AdamW as an optimizer, with parameters set as it is done in Stable Audio Open.
In inference, we use the RMS envelopes predicted though the the video model; after $\mu$-law decoding [^5] https://librosa.org/doc/main/generated/librosa.mu_expand.html, we use interpolation to match the sample rate and employ these envelopes as input to the ControlNet of the audio synthesis model. The model then performs 150 sampling steps to generate the final output, using classifier-free guidance [65] with a scale set to 2.
Regarding the experiments on our Walking The Maps dataset, we fine-tuned our model starting from the best checkpoints obtained after training on Greatest Hits for both the video model and the audio synthesis model. The use of 2 second-long chunks implies a change in the temporal dimensions of the data in our model. Other than that, the parameters used for these experiments are the same as those used for Greatest Hits.
In this case, the video model is fine-tuned with a decreasing learning rate scheduler, with $\gamma=0.5$, for 2500 epochs. While the audio synthesis model is fine-tuned without any change with respect to Greatest Hits experiments.